OpenStudy (anonymous):

Find the distance between the two points (-3,6)(-51,20)

OpenStudy (lgbasallote):

use the distance formula \[\text{distance} = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] \[\text{distance} = \sqrt{[-51 - (-3) ]^2 + (20 - 6)^2}\] does that help?

OpenStudy (anonymous):

yes, thank you!